2. (A simulated economy) Let Y,-F(K,,Lt)-AK) L. The discrete time version of the capital accumulation equation is given by Lo-1 and labor grows at a rate of n-0.01. A-1, Ko-0.01 and 6-0.07 a) Assume that s-0.2. Find the level of consumption in the economy in periods 1,2,3,4 and 100. b) Repeat the computations in (a) assuing s 03 consumption over time after s increases to s = 0.33 d) Do the sanne as in part (c) but assuming that s increases to s = 0.5.
Answer #1

This is a basic simulation exercise and can easily be solved using Excel

Part a)

Please follow these steps:

Keep column A for time period t

Column B for constant A

Column C for variable Kt

Column C for variable Kt

Column D for variable Kt1/3

Column E for variable Lt

Column F for variable Lt2/3

Column G for variable Yt

Formulae in row 2 are as follows:

A2 = 0

B2 = 1

C2 = K0 = 0.01

D2 = K01/3 = 0.01^(1/3)

E2 = L0 = 1

F2 = L02/3= 1^(2/3)

G2 = Y0 = B2*D2*F2

t A K K^(1/3) L L^(2/3) Y
0 1 0.01 0.215443 1 1 0.215443

Formulae for row 3 are as follows:

Formula in row 2 should be as follows:

A3 = t = 1

B3 = A = 1

C3 = K1 =

=((0.2*G2)+((1-0.07)*C2))

where G2 is Y0 and C2 is K0

D3 = K11/3 = C3^(1/3)

E3 = L1 =

=E2*1.01

F3 = L12/3= E3^(2/3)

G3 = Y1= B3*D3*F3

K(1/3) L LA(2/3)Y 1 0.215443 1.01 1.006656 0.052737 0 0.01 0.215443 1 0.052389 0.374179

Copy formulae of row 3 till row 102 which is t = 100

Thus, for s=0.2, level of consumption in the economy is as follows:

Y1 0.377
Y2 0.505
Y3 0.612
Y4 0.705
Y100 4.267

b) Repeat the steps with s = 0.3

C3 = K1 =

=((0.3*G2)+((1-0.07)*C2))

Thus, for s = 0.3, level of consumption in the economy is as follows:

Y1 0.423
Y2 0.588
Y3 0.725
Y4 0.842
Y100 5.226

c) Given that economy is in a steady state and s changes from 0.3 to 0.33, we have to find out what happens to consumption. For this we will do the simulation as conducted in part b for s = 0.33

We see that it reaches steady state at similar time period when s = 0.3 and when s = 0.33 at around t = 36 (% change in consumption becomes 1% Year on year)

s=0.3 s=0.33
Y1 0.423 0.434
Y2 0.588 0.610
Y3 0.725 0.754
Y4 0.842 0.878
Y100 5.226 5.481

d) When s = 0.5

s=0.3 s=0.33 s=0.5
Y1 0.423 0.434 0.492
Y2 0.588 0.610 0.718
Y3 0.725 0.754 0.901
Y4 0.842 0.878 1.057
Y100 5.226 5.481 6.746

As, s increases steady state is reached a little later in time. Moreover, the steady state consumption increases with increase in s.

% change in s s changes from 0.3 to 0.33 = 10% s changes from 0.33 to 0.5 = 52%
Change in Y1 3% 13%
Change in Y2 4% 18%
Change in Y3 4% 19%
Change in Y4 4% 20%
Change in Y100 5% 23%